If the electronegativity difference between two atoms A and B is 2.0, then the percentage distribution character in the molecule is
(A)    54%
(B)    46%
(C)    23%
(D)    72%

Grade:upto college level

2 Answers

Radhika Batra
10 years ago

The percentage ionic character is related to electronegativity difference between the two atoms. It is given by Hannay and Smith relationship.

 

        %age ionic character.

 

                = 16 (XA − XB) + 3.5 (XA − XB)2

 

        Here, XA and XB are elecrtronegativities of two atoms A and B

 

        \ % ionic character = 16(2) + 3.5(2)2 = 46%

 

        or covalent character = 100 – 46 = 54%

naniuppati
10 years ago

b